# A Theory of LLM Sampling: Part Descriptive and Part Prescriptive

### v3

@@ 6 Comparison with human studies
The critical experiment presented in Section 4.1 is inspired by prior work with humans. In Appendix D, we present the results of a similar study conducted on human subjects. We replicate the exact setting using an LLM with a human-like prompt and report the results in a similar visualization to facilitate a direct comparison. Furthermore, we also create exact setting for experiment 3.2 and compare the LLM and human outputs in Appendix E. Finally, the investigation on prototypes presented in Section 5 already follow the same prompt as human studies. Scatter plot of $\hat{\alpha}$ for LLMs and humans show that, though LLMs have strong influence of prescriptive component in sampling due to implicit value associated with each concept, its value system does not correlate with that of humans ($\hat{\alpha}$ pearson correlation of -0.02). The key take away follows that though humans and LLMs have same heuristics in sampling the notion of ideal need not be the same in both. This can lead to undesired manifestation of prescriptive norms in sampling.
@@ 7 Conclusion
In this paper, we set out to better understand the heuristics governing possibility sampling process of LLMs. Based on human cognitive studies, we propose a theory that explains the sampling heuristics to be part descriptive and part prescriptive. However, the exact prescriptive component might not be aligned with humans. As LLMs continue to be integrated into real-world applications, understanding their decision-making heuristics becomes increasingly important. Our results provide a foundational framework for evaluating how LLMs balance statistically probable outcomes with norms of ideality, raising interesting questions about their underlying mechanisms. As a final remark, we would like to emphasize that we do not intend to contribute to “humanizing” AI/ML/LLMs in the way we use terminology or models. Our contribution is intended to draw parallels in behaviour and perform evaluations, as our findings can have an impact on downstream tasks.
@@ 8 Limitations
Although we identify a prescriptive component influencing LLM outputs, the origin of these norms—whether they stem from the pre-training data, reinforcement learning from human feedback (RLHF), or some other aspect of model training—remains under-explored. Further analysis is required to disentangle the contributions of training data versus fine-tuning techniques in shaping prescriptive tendencies in LLMs. Clarifying these origins could inform strategies to better control or mitigate unintended prescriptive biases in model outputs.
Furthermore, this work evaluates prototypicality in LLM similar to evaluation in human subjects. But, prototypicality in neural networks can be studied more closely using their representations. Though the prototypical analysis is stated as an initial exploration in the manuscript, it calls for further research in mechanistic analysis of how prototypes contain prescriptive norms and the possibility of steering and controlling of the prescriptive norm in concept representations.
@@ 9 Ethics and Risks
Furthermore, we hypothesise that this prescriptive norm acts as a foundational bias in other biases found in llms like gender, demography, etc which could be looked at through the lens of value. Since the notion of ideal as shown in the paper is not often aligned with human notions of ideal, different prescriptive components can manifest as different biases under different concepts/domains. This raises important ethical concerns, potentially leading to outputs that do not reflect (a) real-world norms or (b) diverse perspectives. Addressing influence of prescriptive norms is essential for developing transparent, reliable, and just AI technologies, ensuring they contribute positively and ethically across various societal applications.
@@ Appendix E Sampling in relation to existing concepts in humans
In this section, we present the experiment 4.2 on the same concepts and using the same prompt as in prior work in humans by Bear et al. (2020). The results for LLM are shown in Table 5 and the results for humans in the same concepts are shown in Table 4. Comparing this result with the human studies, as shown in Appendix E, we observe that the LLM often gives a ‘strictly ideal’ value when queried for $I(C)$. That is, when a similar question is asked to human test subjects, the number of concepts for which the ideal value is zero is only one. On the other hand, the LLM gives zero for $I(C)$ for 19 concepts (nearly half the time). For instance, the human gives the ideal percentage of ‘high school students underage drinking’ as 13.71%, while the LLM gives $I(C)$ as zero for this concept, showing LLMs, for a lot of concepts, have a notion of stricter ideality compared to the more noisy ideal ratings we seem to observe across humans. We also repeat this experiment for temperature zero as shown in Table 10 in Section K, and observe similar results. We get the following results with other LLMs with default temperatures: Llama-3-7b (binomial $p$ = 0.003), Mixtral-8x7B (binomial $p$ = 0.05), GPT3.5-turbo (binomial $p$ < 0.001), Claude (binomial $p$ < 0.001), Mistral (binomial $p$ = 0.0019).
To illustrate this discrepancy, as shown in figure 6, we present a scatter plot of the $\hat{\alpha}$ values for LLMs and humans. We can see that although the LLM has a strong prescriptive component based on its implicit value associated with each concept, its value system does not correlate with that of humans (Pearson correlation of -0.02). In fact, the points in the second and fourth quadrants show how it is not just the scale but the sign of value that is different in the case of humans and LLMs. This makes the study of prescriptive norms in LLMs more significant as they might not align with human value systems more often than they align. Comparing $\hat{\alpha}$ of humans and the LLM for experiment 5 shows a higher alignment in the value in Figure 6. Here the Pearson correlation of $\hat{\alpha}_{\textit{human}}$ and $\hat{\alpha}_{\textit{LLM}}$ is 0.33. Though not fully aligned in many concepts, only two concepts have different polarities for $\hat{\alpha}$.
@@
Figure 6: Comparing human and LLM on the prototype experiment and sampling on existing concepts. Figure on the left compares from results in Experiment 2 showing some misalignment between LLM and human results due to differences arising in the prescriptive component. Figure on the right compares LLM and human results from Experiment 3 showing more correlation in prototypical concept ratings
@@ Appendix F Motivation for evaluating prototypes
Consider category 4 Exemplar 6 of Grandmother “A 55-year-old woman who likes to party a lot and go out with her friends to casinos and rock concerts. Enjoys playing sports with her grandchildren" (Appendix  P). This example of a grandmother has a lower ideal rating of 5.50 compared to other examples of the category. This is also reflected in the relative lower value of composite example rating (4.5), illustrating that non traditional prototypes are seen less ideally. Similar examples can been be seen in the table in Appendix O.
This implicit bias and punishing of non traditional prototypes has severe implications on tasks where LLM is asked to pick candidates whether it be for academic admissions or hiring processes. Another aspect this bias plays out is between the Exemplar 1 and Exemplar 2 of the Grandmother categrory. Even though Exemplar 2 has lesser average rating compared to Exemplar 1, having a more ideal rating makes it a better example of a grandmother compared to Exemplar 2 illustrating LLMs notion of concepts has a prescriptive norm component.
@@ H.3 Showing effect with different concepts
Appendix J Experiment 2 Topics and Their Sample Questions
@@
Appendix K Experiment two results with temperature zero
@@
Appendix L Experiment two list of prompts
@@
IDEAL NUMBER OF SERVINGS OF FRUITS AND VEGETABLES FOR A PERSON TO CONSUME IN A MONTH ,
@@
IDEAL NUMBER OF MINUTES FOR A PERSON TO SPEND WAITING ON THE PHONE FOR CUSTOMER SERVICE ,
@@
Appendix M Case Study - Patient Recovery time
@@
Appendix N Experiment 1 Glubbing experiment with other LLMs
@@
Appendix O Experiment 3: List of prompts
@@
Appendix P Experiment 3 complete results
@@
Appendix Q Full List of concepts

### v4

@@ 6 Comparison with human studies
The critical experiment presented in Section 4.1 is inspired by prior work with humans Bear et al. (2020). In Appendix D, we present the results of the study conducted on human subjects. We replicate the exact setting using an LLM with a human-like prompts (Appendix J) and report the results in a similar visualization to facilitate a direct comparison. Our results show that, given a new fictional concept with prescriptive and descriptive statistics, humans and LLMs capture these norms and use them to sample options. Morever, we also note that LLMs, like humans seem to have an asymmetric treatment of gains and losses. The undersampling of negative value scenarios is observed to be more than oversampling of the positive value scenarios, potentially pointing to a shared optimism bias. This asymmetry, where both systems avoid negatives more strongly than they pursue positives, can be explored in future work.
Furthermore, we also create exact setting for experiment 3.2 and compare the LLM and human outputs in Appendix E for known concepts. Here we use the same forty concepts used in humans studies to compare the results of LLM and humans. The comparison shows consistency of prescriptive influences across both human cognitive processes and LLM sampling. While both systems exhibit prescriptive components in sampling, a key divergence emerges in their treatment of ideals. Humans consistently conceptualize ideal as modest improvements over statistics (e.g., ideal ‘number of sugary drinks per week’ is $2.41$ while the average is $9.17$), whereas LLMs frequently default to absolute and stricter ideals (e.g., 00 for sugary drinks and $18$ other concepts) indicating moral absolutism which could be a topic of future investigation.
Finally, the investigation on prototypes presented in Section 5 follows the same prompt as human studies. Hence we do not need an explicit recreation of the human prompt for this experiment. This shows initial evidence that concept prototypicality scoring in LLMs are driven by the same components as in humans. Interestingly, scatter plot of $\hat{\alpha}$ for LLMs and humans for prototypicality 6 (right) show that, the amplitude of the influence of prescriptive component in sampling also correlate with that of humans ($\hat{\alpha}$ Pearson correlation of 0.33).
As a final note, our experiments probe systematic patterns in LLM outputs, revealing biases and decision tendencies analogous to human behavioral studies. We do not presuppose that these patterns emerge from human-like reasoning mechanisms.
@@ 7 Conclusion
In this paper, we set out to better understand the heuristics governing possibility sampling process of LLMs. Based on human cognitive studies, we propose a theory that explains the sampling heuristics to be part descriptive and part prescriptive. However, the exact prescriptive component might not be aligned with humans. As LLMs continue to be integrated into real-world applications, understanding their decision-making heuristics becomes increasingly important. Our results provide a foundational framework for evaluating how LLMs balance statistically probable outcomes with norms of ideality, raising interesting questions about their underlying representations. As a final remark, we would like to emphasize that we do not intend to contribute to “humanizing” AI/ML/LLMs in the way we use terminology or models. Instead, our contribution is intended to draw parallels in behaviour and perform evaluations, as our findings can have an impact on downstream tasks.
8 Acknowledgements
This work was partially funded by ELSA – European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No.101070617. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them. The project on which this report is based was funded by the Federal Ministry of Education and Research under the funding code 16KIS2012. The responsibility for the content of this publication lies with the authors.
@@ 9 Limitations
Although we identify a prescriptive component influencing LLM outputs, the origin of these norms, whether they stem from the pre-training data, reinforcement learning from human feedback (RLHF), or some other aspect of model training remains under-explored. Further analysis is required to disentangle the contributions of training data versus fine-tuning techniques in shaping prescriptive tendencies in LLMs. Clarifying these origins could inform strategies to better control or mitigate unintended prescriptive biases in model outputs. The paper also does not explore the mechanism of how norms affect heuristics.
Furthermore, this work evaluates prototypicality in LLM similar to evaluation in human subjects. But, prototypicality in neural networks can be studied more closely using their representations. Though the prototypical analysis is stated as an initial exploration in the manuscript, it calls for further research in mechanistic analysis of how prototypes contain prescriptive norms and the possibility of steering and controlling of these norms in concept representations.
@@ 10 Ethics and Risks
Furthermore, we hypothesise that this prescriptive norm acts as a foundational bias in other biases found in LLMs like gender, demography, etc which could be looked at through the lens of value. Since there are no guarantees on the ideals of LLMs, LLM sampled options can appear with different biases under different concepts/domains. This raises important ethical concerns, potentially leading to outputs that do not reflect (a) real-world norms or (b) diverse perspectives. Addressing influence of prescriptive norms is essential for developing transparent, reliable, and fair AI technologies, ensuring they contribute positively and ethically across various societal applications.
@@ Appendix E Sampling in relation to existing concepts in humans
In this section, we present the experiment 4.2 on the same concepts and using the same prompt as in prior work in humans by Bear et al. (2020). The results for LLM are shown in Table 5 and the results for humans in the same concepts are shown in Table 4. Comparing this result with the human studies, as shown in Appendix E, we observe that the LLM often gives a ‘strictly ideal’ value when queried for $I(C)$. That is, when a similar question is asked to human test subjects, the number of concepts for which the ideal value is zero is only one. On the other hand, the LLM gives zero for $I(C)$ for 19 concepts (nearly half the time). For instance, the human gives the ideal percentage of ‘high school students underage drinking’ as 13.71%, while the LLM gives $I(C)$ as zero for this concept, showing LLMs, for a lot of concepts, have a notion of stricter ideality compared to the more noisy ideal ratings we seem to observe across humans.
We also repeat this experiment for temperature zero as shown in Table 10 in Section L, and observe similar results. We get the following results with other LLMs with default temperatures: Llama-3-7b (binomial $p$ = 0.003), Mixtral-8x7B (binomial $p$ = 0.05), GPT3.5-turbo (binomial $p$ < 0.001), Claude (binomial $p$ < 0.001), Mistral (binomial $p$ = 0.0019).
We present a scatter plot of $\hat{\alpha}$ values for LLMs and humans in Figure 6. We see that although the LLM has a strong prescriptive component based on its implicit value associated with each concept, its amplitude of shift towards prescriptive norm does not correlate with that of humans (Pearson correlation of -0.02). This makes the study of prescriptive norms in LLMs more significant as they might not manifest in samples like it does in humans. Comparing $\hat{\alpha}$ of humans and the LLM for experiment 5 shows a higher alignment in the shift (Figure 6). Here the Pearson correlation of $\hat{\alpha}_{\textit{human}}$ and $\hat{\alpha}_{\textit{LLM}}$ is 0.33.
@@
Figure 6: Comparing human and LLM on the prototype experiment and sampling on existing concepts. Figure on the left compares from results in Experiment 2 showing some misalignment between LLM and human results due to differences arising in the prescriptive component. Figure on the right compares LLM and human results from Experiment 3 showing more correlation in prototypical concept ratings.
@@ Appendix F Motivation for evaluating prototypes
Consider category 4 Exemplar 6 of Grandmother “A 55-year-old woman who likes to party a lot and go out with her friends to casinos and rock concerts. Enjoys playing sports with her grandchildren" (Appendix  Q). This example of a grandmother has a lower ideal rating of 5.50 compared to other examples of the category. This is also reflected in the relative lower value of composite example rating (4.5), illustrating that non traditional prototypes are seen less ideally. Similar examples can been be seen in the table in Appendix P.
This implicit bias and punishing of non traditional prototypes has severe implications on tasks where LLM is asked to pick candidates whether it be for academic admissions or hiring processes. Another aspect this bias plays out is between the Exemplar 1 and Exemplar 2 of the Grandmother category. Even though Exemplar 2 has lesser average rating compared to Exemplar 1, having a more ideal rating makes it a better example of a grandmother compared to Exemplar 2 illustrating LLMs notion of concepts has a prescriptive norm component.
@@ H.3 Showing effect with different concepts
Prompts of Experiment 1 - Glubbing (Positive Case) Suppose there is a hobby called glubbing. Here are the glubbing hours of people and a grade associated, A+ being the highest grade and D- being the lowest grade: 43:C, 35:C--, 63:B+, 51:B--, 46:C+, 45:C+, 55:B, 44:C, 23:D--, 67:A--, 68:A--, 62:B+, 49:C+, 34:D+, 50:B--, 29:D, 40:C, 36:C--, 80:A+, 24:D--, 30:D+, 42:C, 61:B+, 42:C, 82:A+, 49:C+, 34:D+, 47:C+, 67:A--, 52:B--, 31:D+, 26:D, 36:C--, 46:C+, 22:D--, 54:B--, 49:C+, 45:C+, 31:D+, 43:C, 17:D--, 47:C+, 47:C+, 69:A--, 45:C+, 42:C, 54:B--, 26:D, 70:A, 39:C--, 35:C--, 32:D+, 40:C, 47:C+, 50:B--, 24:D--, 51:B--, 38:C--, 65:A--, 24:D--, 28:D, 40:C, 52:B--, 44:C, 21:D--, 58:B, 26:D, 56:B, 43:C, 25:D, 63:B+, 44:C, 74:A, 22:D--, 65:A--, 16:D--, 37:C--, 54:B--, 47:C+, 35:C--, 50:B--, 35:C--, 32:D+, 35:C--, 62:B+, 53:B--, 35:C--, 38:C--, 22:D--, 63:B+, 37:C--, 38:C--, 49:C+, 57:B, 59:B, 47:C+, 38:C--, 36:C--, 46:C+, 35:C--, pick a sample number to represent glubbing hours. The value should be an integer value between 0 and 100. Print only the number and not the complete sentence.
Prompts of Experiment 1 - Glubbing- Negative Suppose there is a hobby called glubbing. Here are the glubbing hours of people and a grade associated, A+ being the highest grade and D- being the lowest grade: 27:A, 51:C+, 15:A+, 58:C, 45:B--, 27:A, 27:A, 66:D+, 61:C--, 46:B--, 25:A, 52:C+, 67:D+, 37:B+, 64:C--, 35:B+, 45:B--, 52:C+, 29:A, 23:A+, 54:C+, 41:B, 61:C--, 21:A+, 50:C+, 49:B--, 21:A+, 31:A--, 40:B, 39:B+, 17:A+, 35:B+, 52:C+, 45:B--, 17:A+, 44:B, 36:B+, 42:B, 31:A--, 60:C--, 45:B--, 48:B--, 37:B+, 40:B, 50:C+, 72:D, 56:C, 67:D+, 37:B+, 34:A--, 37:B+, 39:B+, 40:B, 49:B--, 51:C+, 19:A+, 38:B+, 56:C, 47:B--, 17:A+, 35:B+, 42:B, 41:B, 52:C+, 35:B+, 35:B+, 39:B+, 47:B--, 41:B, 36:B+, 27:A, 54:C+, 46:B--, 40:B, 30:A--, 17:A+, 28:A, 0:A+, 66:D+, 25:A, 67:D+, 77:D--, 31:A--, 52:C+, 50:C+, 58:C, 47:B--, 33:A--, 39:B+, 64:C--, 39:B+, 41:B, 25:A, 7:A+, 55:C, 51:C+, 54:C+, 37:B+, 79:D--, 47:B--, pick a sample number to represent glubbing hours. The value should be an integer value between 0 and 100. Print only the number and not the complete sentence.
Appendix J Experiment 6 for human comparison
The full prompts for exact comparison of LLM results with humans. Reported for all 3 cases of the experiment ,the positive ideal, negative ideal and neutral respectively.
@@
Appendix K Experiment 2 Topics and Their Sample Questions
@@
Appendix L Experiment two results with temperature zero
@@
Appendix M Experiment two list of prompts
@@
IDEAL NUMBER OF SERVINGS OF FRUITS AND VEGETABLES FOR A PERSON
TO CONSUME IN A MONTH ,
@@
IDEAL NUMBER OF MINUTES FOR A PERSON TO SPEND WAITING ON
THE PHONE FOR CUSTOMER SERVICE ,
@@
Appendix N Case Study - Patient Recovery time
@@
Appendix O Experiment 1 Glubbing experiment with other LLMs
@@
Appendix P Experiment 3: List of prompts
@@
Appendix Q Experiment 3 complete results
@@
Appendix R Full List of concepts